OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form. (-5,-4) and (15,0)

OpenStudy (anonymous):

It is simple first take out slope m=(y2-y1)/(x2-x1) m=(15+5)/(0+4) =20/4 =5 m=5 use this equation since you do not have the value of c (y-y1)=m(x-x1) I am taking (-5,-4) as my co-ordinates so my x1=-5 and y2=-4 m=5 y+4=5(x+5) y+4=5x+20 y=5x+20-4 y=5x+16 therefore slope intercept form is y=5x+16 with m=5 and c=16 (y-intercept)